# Accelerated Gradient Methods with Biased Gradient Estimates: Risk Sensitivity, High-Probability Guarantees, and Large Deviation Bounds

### v1

@@ Appendix A Existing bounds on the $H_{\infty}$-norm for $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$
$\quad\times\exp\Bigg{\{}\bar{\theta}\Bigg{[}c_{k+1,K}(\bar{\theta})+\frac{1}{\bar{\theta}}\log\mathbb{E}\left[e^{\frac{\tilde{t}_{k+1,K}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k+1}}}\|^{2}}{2\sigma^{2}}}\,\Big{|}\,\xi_{k}=\xi^{+},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{k-1}=\xi}\right]\Bigg{]}\Bigg{\}}$
@@
$\quad\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\exp\Bigg{\{}\bar{\theta}\Bigg{[}c_{k+1,K}(\bar{\theta})+\frac{1}{\bar{\theta}}\log\left(\frac{1+\tilde{t}_{k+1,K}}{1-\tilde{t}_{k+1,K}}\right)\Bigg{]}\Bigg{\}}}\,$,
@@ Appendix E Supporting Lemmas
$\phi(\theta,\check{a},\check{b})=\begin{cases}\check{a}+4\sigma^{2}\hat{b}&\mbox{if}\quad\theta=0,\\ \check{a}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\frac{2\sigma^{2}}{\theta}\log\left(\frac{1+\theta\check{b}}{1-\theta\check{b}}\right)&\mbox{if}\quad 0<\theta<{1}/{\check{b}},\\ +\infty&\mbox{if}\quad\theta\geq{1}/{\check{b}}\mbox{ or }\theta<0,\end{cases}$
@@
is non-increasing, i.e. $\hat{\phi}_{j_{1}}(\theta)\geq\hat{\phi}_{j_{2}}(\theta)$ for every $\theta$ whenever $0\leq j_{1}\leq j_{2}\leq K-1$. Therefore, in the setting of theorem˜5.1, we have
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\sup_{0\leq\theta<\frac{1}{(\overline{H}_{\infty})^{2}}}\frac{\theta}{2\sigma^{2}}\left(t-\phi(\theta,\check{a}_{K}^{(2)},\check{b}_{K}^{(2)})\right)\bigg{\}}$

### v2

@@ Appendix A Existing bounds on the $H_{\infty}$-norm for $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$
$\qquad\times\exp\Bigg\{\bar{\theta}\Bigg[c_{k+1,K}(\bar{\theta})+\frac{1}{\bar{\theta}}\log\mathbb{E}\left[e^{\frac{\tilde{t}_{k+1,K}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k+1}}}\|^{2}}{2\sigma^{2}}}\,\Big|\,\xi_{k}=\xi^{+},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\xi_{k-1}=\xi}\right]\Bigg]\Bigg\}$
@@
$\qquad\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\exp\Bigg\{\bar{\theta}\Bigg[c_{k+1,K}(\bar{\theta})+\frac{1}{\bar{\theta}}\log\left(\frac{1+\tilde{t}_{k+1,K}}{1-\tilde{t}_{k+1,K}}\right)\Bigg]\Bigg\}}\,$,
@@ Appendix E Supporting Lemmas
$\phi(\theta,\check{a},\check{b})=\begin{cases}\check{a}+4\sigma^{2}\check{b}&\mbox{if}\quad\theta=0,\\ \check{a}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\frac{2\sigma^{2}}{\theta}\log\left(\frac{1+\theta\check{b}}{1-\theta\check{b}}\right)&\mbox{if}\quad 0<\theta<{1}/{\check{b}},\\ +\infty&\mbox{if}\quad\theta\geq{1}/{\check{b}}\mbox{ or }\theta<0,\end{cases}$
@@
is non-increasing, i.e., $\hat{\phi}_{j_{1}}(\theta)\geq\hat{\phi}_{j_{2}}(\theta)$ for every $\theta$ whenever $0\leq j_{1}\leq j_{2}\leq K-1$. Therefore, in the setting of theorem˜5.1, we have
@@
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\sup_{0\leq\theta<\frac{1}{(\overline{H}_{\infty})^{2}}}\frac{\theta}{2\sigma^{2}}\left(t-\phi(\theta,\check{a}_{K}^{(2)},\check{b}_{K}^{(2)})\right)\bigg\}$